OpenStudy (anonymous):

What is the sum of 81 + i squared? The rule is that i squared is -1 . the then is 81 + i squared = 81 + (-1) which 81-1? .

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

\(\bf i^2\implies \sqrt{-1}\cdot \sqrt{-1}\implies (\sqrt{-1})^2\implies \sqrt{(-1)^2}\implies -1\)